OpenStudy (anonymous):

Give me 3 coordnates that form a triangle. Also give me the slope and y intercept of these coordnates.

OpenStudy (anonymous):

Please help

OpenStudy (lalaly):

A(3,5), B(4,7), C(7,6)

OpenStudy (anonymous):

Slope , y-intercept?

OpenStudy (lalaly):

slope of AB=7-5/4-3 = 2 slope of AC=6-5/7-3 = 1/4 cntinue

OpenStudy (anonymous):

How can u have the y-intercept of a co-ordinate?